OpenStudy (anonymous):

simplify (-5+6i)(2+5i)

OpenStudy (anonymous):

Use FOIL to start with. Then remember that \(i^2 = -1\).

OpenStudy (anonymous):

-10-13i?

OpenStudy (anonymous):

No. That's not what I get. Did you FOIL it first?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok lets see

OpenStudy (anonymous):

Can you show me your FOIL step?

OpenStudy (anonymous):

ok first your answer is not correct

OpenStudy (neer2890):

\[(-5+6i)(2+5i)=-5*2+(-5)*5i+6i*2+6i*5i=-10-25i+12i+30i ^{2}\] \[=-10-13i+30i ^{2}\] and we know that \[i ^{2}=-1\] so \[=-10-13i+30i ^{2}=-10-13i+30*(-1)=-10-13i-30=-40-13i\]